# Preconditioned block encodings for quantum linear systems

### v1

@@ Abstract
Quantum linear system solvers like the Quantum Singular Value Transformation (QSVT) require a block encoding of the system matrix $A$ within a unitary operator $U_{A}$. Unfortunately, block encoding often results in significant subnormalisation and increase in the matrix’s effective condition number $\kappa$, affecting the efficiency of solvers. Matrix preconditioning is a well-established classical technique to reduce $\kappa$ by multiplying $A$ by a preconditioner $P$. Here, we study quantum preconditioning for block encodings. We consider four preconditioners and two encoding approaches: (a) separately encoding $A$ and its preconditioner $P$, followed by quantum multiplication, and (b) classically multiplying $A$ and $P$ before encoding the product in $U_{PA}$. Their impact on subnormalisation factors and condition number $\kappa$ are analysed using practical matrices from Computational Fluid Dynamics (CFD). Our results show that (a) quantum multiplication introduces excessive subnormalisation factors, negating improvements in $\kappa$. We introduce preamplified quantum multiplication to reduce subnormalisation, which is of independent interest. Conversely, we see that (b) encoding of the classical product can significantly improve the effective condition number using the Sparse Approximate Inverse preconditioner with infill. Further, we introduce a new matrix filtering technique that reduces the circuit depth without adversely affecting the matrix solution. We apply these methods to reduce the number of QSVT phase factors by a factor of 25 for an example CFD matrix of size 1024x1024.
@@ 1.1 Previous work
A quantum algorithm for solving a circulant matrix system using a modified HHL solver [13] demonstrated the benefit of being able to express circulant matrices in terms of Quantum Fourier Transforms. Separately, [14] showed how to implement a circulant preconditioner using a modified quantum singular value estimation scheme.
@@
A fast inversion scheme for preconditioning [16] writes the matrix system as $(A+B)\ket{x}=\ket{b}$ and assumes $||A||>>||B||$. Rewriting gives:
$(I+A^{-1}B)\ket{x}=A^{-1}\ket{b}$ (3)
$A$ is fast-invertible if, after rescaling $A$ so that $||A^{-1}||=1$, there is a block-encoding of $A^{-1}$ where the number of oracle queries for $A$ is not dependent on its condition number. Diagonal and 1-sparse matrices can be fast-inverted. Although for general matrices, the scheme requires a classical Singular Value Decomposition to give $A=UDV^{\dagger}$. For normal matrices, $U=V$ and in certain cases $V$ is easily computed, e.g. when $V$ is the Quantum Fourier Transform. A comparison of the SPAI, circulant and fast inversion preconditioners [17] showed significant scaling benefits for fast inversion using an inverse Laplacian preconditioner. The test case focussed on fracture mechanics and split the matrix into $A=\Delta+A_{F}$, where $\Delta$ is the Laplacian of the flow field in the absence of fractures and $A_{F}$ is the fracture matrix. For the cases studied, the Laplacian has a known eigenvalue decomposition. Classical preconditioning using an inverse Laplacian [18, 19, 20, 21] has demonstrated similar benefits using Krylov subspace iterative solvers.
Whilst previous work has provided some practical details on preconditioning quantum linear solvers, none has considered the impact of the subnormalisation factors arising from block encoding. Nor has the difference between the circuit implementations of classical and quantum multiplication of $P$ and $A$ been considered.
@@ 1.2 This work
Next, in Section 3 we study preconditioning. Many classical methods do not explicitly form the matrices $A$ and $P$, instead subroutines are called to perform the matrix-vector multiplication. For QLES, the classical overhead of forming $PA$ suggests the analogous approach of multiplying the encoded matrices for $A$ and $P$ rather than encoding $PA$. For compactness, we use the terms quantum multiplication for the separate encoding of $P$ and $A$ and classical multiplication for pre-multiplication of $P$ and $A$ prior to encoding the product $PA$. We introduce preamplified quantum multiplication (Section 3.3.1), which improves the state of the art for multiplying block encodings and is of independent interest. Both approaches (classical multiplication and quantum multiplication) are studied and compared: We investigate how the subnormalisation factors associated with the block-encoding of $A$, $P$ and $PA$ affect the condition number for each of the preconditioners considered.
@@
Finally, in Section 5 we turn to CFD applications of the methodology presented. Using emulated circuits for CFD matrices ranging from $16\times 16$ to $096\times 4,$ 096 × 4 , 096 we provide the first practical circuit level evaluations of preconditioning with QSVT.
@@ 2.2 Diagonal scaling and encoding the original matrix
Since $A$ is a banded diagonal matrix, it uses the encoding method described in Section 2.4. However, we first describe the encoding of Toeplitz matrices as the banded diagonal encoding is a generalisation thereof.
@@ 2.3 Toeplitz matrix encoding
Figure 1: Circuit for encoding a 16x16 pentadiagonal Toeplitz matrix with diagonals at offsets: $-4$, $-1$, 0, 1, 4. Negative offsets are super-diagonals, positive offsets are sub-diagonals and 0 is the main diagonal. Strictly, the above circuit has a controlled addition of 0, but this is omitted as it has no effect.
Figure 1 gives an overview of the circuit to encode a 16x16 pentadiagonal Toeplitz matrix based on a 4x4 CFD pressure correction equation. Each diagonal has a constant value; these are loaded using the prep operator. The multiplexed controls then offset each diagonal by a positive amount. Note that there is no offset for the $-4$ diagonal as a controlled addition of 0 has no effect. After the controlled additions, the final arithmetic operator moves the diagonals to their correct location.
The qubit labels follow [22] where those labelled |$\ket{j}$ designate the columns of the matrix, and those labelled |$\ket{s}$ designate the diagonals. The |$\ket{del}$ qubit is needed so that the encoding creates a 32x32 matrix. This is because when a diagonal is offset, values that move out of the range of the matrix wrap round to the other side. This contaminates the 32x32 matrix but not the upper left 16x16 block.
@@ 2.4 Banded diagonal matrix encoding
Figure 2 shows a schematic of the circuit to encode a 16x16 pentadiagonal matrix. As before, the qubit labels follow [22] where the data loading controlled rotations are applied to the |$\ket{d_{0}}$ qubit. The encoding oracle follows full matrix encoding [24] but instead of using row-column indexing, diagonal-column indexing is used. Importantly, the subnormalisation factor scales with the number of diagonals rather than with the dimension of the matrix. The operators $D_{0},D_{-1},...$ contain both the data loading oracles and the arithmetic operators to offset the diagonals to the correct position. A negative offset creates a super-diagonal and a positive offset creates a sub-diagonal.
@@
Figure 3 shows the circuit segment for loading a banded tridiagonal matrix. The controls on the |$\ket{j}$ qubits select the column and the controls on the |$\ket{s}$ qubits select the diagonal. By themselves, the controlled rotations create a direct sum of the three diagonal matrices. The controlled subtraction offsets the part of the direct sum containing the super-diagonal by 1. Similarly, the controlled addition offsets the sub-diagonal by 4. The prep-unprep operations scale and mix the diagonals to give the full encoded matrix in the top left corner of the resulting unitary. The $X$ gate after the rotations is required for arcsin encoding [25]. Since only the non-zero entries of the matrix are encoded, there is no out of range wrap-around as occurs in the Toeplitz encoding.
@@ 2.5 Circulant matrix encoding
From Equation 36 in Appendix D, the circulant preconditioner must load $F\Lambda^{-1}F^{\dagger}$ where $F$ is the Fourier transform. The diagonal matrix of inverse eigenvalues, $\Lambda^{-1}$, can be computed from Equation 35 and can be encoded using diagonal encoding. The resulting circuit including the product with $A$ is shown in Figure 4
@@ 2.6 Matrix infill
Although the inverse of a sparse matrix is not, itself, sparse, the default for many preconditioners is to impose the same sparsity pattern on $P$ as the original matrix $A$. This limits the effectiveness of the preconditioning but is often necessary due to computational limits. An alternative is for the preconditioner to use infill where $P$ is allowed to have more non-zeros than $A$. This is discussed in Appendix B and illustrated in Figure 12. For banded diagonal matrices, the infill procedure introduces additional diagonals. The infill algorithm can be repeated with each repetition giving a better approximation to $A^{-1}$ at the expense of adding more diagonals.
@@
Infill levels diag($P$) diag($PA$) diag($P$) diag($PA$) non-zero diag($PA$)
@@
Table 1: Number of banded diagonals in $P$ and the classical product $PA$ for different infill levels with the TPAI and SPAI preconditioners. The final column shows the number of diagonals that have non-zero entries.
Table 1 shows the number of diagonals with infill levels from zero to three for $P$ and the classical product $PA$ using the TPAI and SPAI preconditioners. Note that CLAI generates a dense matrix for which infill is not relevant. As described in Appendix B and Appendix C, the preconditioners use different infill strategies with SPAI adding more diagonals per infill operation.
An unexpected feature of SPAI is that a number of diagonals in the product $PA$ consist entirely of zeros to within the precision of the arithmetic. The is a result of exactly solving each reduced system as described in Section B.1. For SPAI with three levels of infill, performing quantum multiplication requires encoding two banded matrices with 5 and 41 diagonals. Whereas, classical multiplication encodes a single banded matrix with 21 diagonals. Of course, the latter has a high classical pre-processing cost to compute $PA$ prior to encoding.
@@ 3 Preconditioning
As a test case for preconditioning, we use CFD pressure correction matrices for the flow in a lid-driven cavity taken from the open source qc-cfd  [^1] https://github.com/rolls-royce/qc-cfd/tree/main/2D-Cavity-Matrices repository. The CFD meshes range from $4\times 4$ to $64\times 64$ cells. The corresponding pressure correction matrices have dimensions ranging from $16\times 16$ to $096\times 4,$ 096 × 4 , 096 and are each extracted from the non-linear solver after 100 outer iterations.
@@
We study the impact of preconditioning on subnormalisation and subnormalised condition number, where Table 2 gives an overview of the scenarios considered. Firstly, the preconditioners are evaluated exactly as they would be used by a classical solver (Section 3.1). Then we study the subnormalisation factors for each preconditioner (Section 3.2), required for preconditioning by quantum multiplication. In Sections 3.3 and 3.4, we study preconditioning with the quantum multiplication and classical multiplication approaches, respectively. For preconditioning by classical multiplication, only TPAI and SPAI are considered as the CLAI preconditioner produces a dense matrix that would require excessive classical computing resources. The effects of preconditioning on both subnormalisation factors and condition numbers are considered.
As mentioned above diagonal scaling is a preprocessing method used by all the preconditioners. It is included in the classical section only to show the small effect it has. The SPAI and TPAI results include infill up to three levels. All matrices are scaled to have $||A||_{max}=1$ prior to block encoding.
@@ 3.1 Classical preconditioning
Figure 5: Classical preconditioning - condition numbers for $PA$ as used by a classical algorithm. Computed using simple matrix-matrix multiplication. Numbers in brackets indicate levels of infill.
Figure 5 shows the effect of the four preconditioners on the condition numbers of the CFD matrix. As anticipated, diagonal scaling has little benefit by itself. There is some erratic behaviour for SPAI and TPAI with higher levels of infill on the coarser meshes which is due to the small matrix size. For larger meshes, there is an increasing benefit of more levels of infill. TPAI shows an initial improvement without infill but only a marginal improvement thereafter. There is no option to use infill with CLAI. However, it performs as well as SPAI with 3 levels of infill for the larger matrices.
@@ 3.2 Subnormalisation factors of preconditioners
We first consider the subnormalisation factors for encoding the preconditioner $P$. From Figure 6(a) we can see that the subnormalisation factor for $A$ is approximately $s=2$ for all meshes. This is due to the mass conservation equation being used to construct the pressure correction matrix, where for each row the diagonal entry is the negative of the sum of the off diagonal entries. The small difference from $2$ is due to the fact that each diagonal is scaled by $1/||D||_{max}$ prior to encoding. The locations of the maximal entries can occur on different rows of the matrix and, hence, $s=2$ is lower bound if all the maxima occur on the same row. However, this is small effect that reduces on the finer meshes.
(a) Diagonal preconditioning, $s$.
@@
The first observation is the striking difference between the values of $s$ for the SPAI and TPAI preconditioners, Figure 6(b) and Figure 6(c), as the level of infill is increased. This is largely due to the difference in the way each preconditioner adds infill. The numbers of diagonals added for each level of infill are compared in Table 1. These are independent of the CFD mesh size except for the two smallest cases. The difference in subnormalisation factors is largely attributable to the difference in the number of diagonals. This is not a pro-rata increase as it is the maximum entry along each diagonal that sets the prep coefficients.
@@
The second observation is that for the circulant preconditioner, $s$ rises rapidly with the size of the matrix. This is due to the fact that as the mesh size increases, the cell dimensions become smaller and the lowest eigenvalue reduces. The increase in $s$ is driven entirely by the need to scale $\Lambda^{-1}$ so that $||\Lambda^{-1}||_{max}=1$. There is no subnormalisation factor associated with the quantum Fourier transforms.
(a) Diagonal preconditioning, $\kappa_{s}$.
@@
Figure 7: Preconditioning by quantum multiplication - condition numbers using all preconditioners. The blue line (Original) gives the classical condition numbers ($\kappa$) of $A$ without preconditioning. The orange line (Diagonal) gives the quantum condition numbers ($\kappa_{s}$) for the encoded matrix $U_{A}$ without preconditioning. Numbers in brackets indicate levels of infill.

### v3

@@ Abstract
Quantum linear system solvers like the Quantum Singular Value Transformation (QSVT) require a block encoding of the system matrix $A$ within a unitary operator $U_{A}$. Unfortunately, block encoding often results in significant subnormalisation and increase in the matrix’s effective condition number $\kappa$, affecting the efficiency of solvers. Matrix preconditioning is a well-established classical technique to reduce $\kappa$ by multiplying $A$ by a preconditioner $P$. Here, we study quantum preconditioning for block encodings. We consider four preconditioners and two encoding approaches: (a) separately encoding $A$ and its preconditioner $P$, followed by quantum multiplication, and (b) classically multiplying $A$ and $P$ before encoding the product in $U_{PA}$. Their impact on subnormalisation factors and condition number $\kappa$ are analysed using practical matrices from Computational Fluid Dynamics (CFD). Our results show that (a) quantum multiplication introduces excessive subnormalisation factors, negating improvements in $\kappa$. We study preamplified quantum multiplication to reduce subnormalisation. Conversely, we see that (b) encoding of the classical product can significantly improve the effective condition number using the Sparse Approximate Inverse preconditioner with infill. Further, we introduce a new matrix filtering technique that reduces the circuit depth without adversely affecting the matrix solution. We apply these methods to reduce the number of QSVT phase factors by a factor of 25 for an example CFD matrix of size 1024x1024.
@@ 1.1 Previous work
A quantum algorithm for solving a circulant matrix system using a modified HHL solver [13] demonstrated the benefit of being able to express circulant matrices in terms of Quantum Fourier Transforms. Separately, [14] showed how to implement a circulant preconditioner using a modified quantum singular value estimation scheme. A circuit implementation for circulant matrices including the multiplication of two circulant matrices was developed by [15]. The encoding followed the prep-select method and the matrix product method followed [6]. The circulant system of equations was solved using HHL.
@@
[…]
AA is fast-invertible if, after rescaling $A$ so that $A^{-1}||=1$, there is a block-encoding of $A^{-1}$ where the number of oracle queries for $A$ is not dependent on its condition number. Diagonal and 1-sparse matrices can be fast-inverted. Although for general matrices, the scheme requires a classical Singular Value Decomposition to give $A=UDV^{\dagger}$. For normal matrices, $U=V$ and in certain cases $V$ is easily computed, e.g. when $V$ is the Quantum Fourier Transform. A comparison of the SPAI, circulant and fast inversion preconditioners [18] showed significant scaling benefits for fast inversion using an inverse Laplacian preconditioner. The test case focussed on fracture mechanics and split the matrix into $A=\Delta+A_{F}$, where $\Delta$ is the Laplacian of the flow field in the absence of fractures and $A_{F}$ is the fracture matrix. For the cases studied, the Laplacian has a known eigenvalue decomposition. Classical preconditioning using an inverse Laplacian [19, 20, 21, 22] has demonstrated similar benefits using Krylov subspace iterative solvers.
The likely effect of subnormalisation factors on quantum preconditioning for solving elliptic PDEs was considered by [23]. Instead of either quantum or classical multiplication, the preconditioned matrix $PA$ was formed directly on the classical computer using the multi-level BPX preconditioner [24].
Whilst previous work has provided some practical details on preconditioning quantum linear solvers, none has explicitly considered the practical impact of the subnormalisation factors arising from block encoding. Nor has the difference between the circuit implementations of classical and quantum multiplication of $P$ and $A$ been considered.
@@ 1.2 This work
Next, in Section 3 we study preconditioning. Many classical methods do not explicitly form the matrices $A$ and $P$, instead subroutines are called to perform the matrix-vector multiplication. For QLES, the classical overhead of forming $PA$ suggests the analogous approach of multiplying the encoded matrices for $A$ and $P$ rather than encoding $PA$. For compactness, we use the terms quantum multiplication for the separate encoding of $P$ and $A$ and classical multiplication for pre-multiplication of $P$ and $A$ prior to encoding the product $PA$. We also consider preamplified quantum multiplication (Section 3.3.1), which can improve the figure of merit when multiplying block encodings. Both approaches (classical multiplication and quantum multiplication) are studied and compared: We investigate how the subnormalisation factors associated with the block-encoding of $A$, $P$ and $PA$ affect the condition number for each of the preconditioners considered.
@@
Finally, in Section 5 we turn to CFD applications of the methodology presented. Using emulated circuits for CFD matrices ranging from $16\times 16$ to $4,096\times 4,096$ we provide the first practical circuit level evaluations of preconditioning with QSVT.
@@ 2.2 Diagonal scaling and encoding the original matrix
Since $A$ is a banded diagonal matrix, it uses the encoding method described in Section˜2.4. However, we first describe the encoding of Toeplitz matrices as the banded diagonal encoding is a generalisation thereof.
@@ 2.3 Toeplitz matrix encoding
Figure 1: Circuit for encoding a 16x16 pentadiagonal Toeplitz matrix with diagonals at offsets: 4-4, 1-1, 0, 1, 4. Negative offsets are super-diagonals, positive offsets are sub-diagonals and 0 is the main diagonal. Strictly, the above circuit has a controlled addition of 0, but this is omitted as it has no effect.
Figure˜1 gives an overview of the circuit to encode a 16x16 pentadiagonal Toeplitz matrix based on a 4x4 CFD pressure correction equation. Each diagonal has a constant value; these are loaded using the prep operator. The multiplexed controls then offset each diagonal by a positive amount. Note that there is no offset for the 4-4 diagonal as a controlled addition of 0 has no effect. After the controlled additions, the final arithmetic operator moves the diagonals to their correct location.
[…]
@@ 2.4 Banded diagonal matrix encoding
[…]
@@
[…]
@@ 2.5 Circulant matrix encoding
From Equation˜36 in Appendix˜D, the circulant preconditioner must load $F\Lambda^{-1}F^{\dagger}$ where $F$ is the Fourier transform. The diagonal matrix of inverse eigenvalues, $\Lambda^{-1}$, can be computed from Equation˜35 and can be encoded using diagonal encoding. The resulting circuit including the product with $A$ is shown in Figure˜4
@@ 2.6 Matrix infill
Although the inverse of a sparse matrix is not, itself, sparse, the default for many preconditioners is to impose the same sparsity pattern on $P$ as the original matrix $A$. This limits the effectiveness of the preconditioning but is often necessary due to computational limits. An alternative is for the preconditioner to use infill where $P$ is allowed to have more non-zeros than $A$. This is discussed in Appendix˜B and illustrated in Figure˜12. For banded diagonal matrices, the infill procedure introduces additional diagonals. The infill algorithm can be repeated with each repetition giving a better approximation to $A^{-1}$ at the expense of adding more diagonals.
@@
Infill levels diag($P$) diag(PAPA) diag($P$) diag(PAPA) non-zero diag($PA$)
@@
$i$ - - $5+2i(i+3)$ $5+2(i+1)(i+4)$ $9+4i$
Table 1: Number of banded diagonals in $P$ and the classical product PAPA for different infill levels with the TPAI and SPAI preconditioners. The final column shows the number of diagonals that have non-zero entries. The final row shows the asymptotic complexity for SPAI.
Table˜1 shows the number of diagonals with infill levels from zero to three for $P$ and the classical product $PA$ using the TPAI and SPAI preconditioners. Note that CLAI generates a dense matrix for which infill is not relevant. As described in Appendix˜B and Appendix˜C, the preconditioners use different infill strategies with SPAI adding more diagonals per infill operation.
An unexpected feature of SPAI is that a number of diagonals in the product $PA$ consist entirely of zeros to within the precision of the arithmetic. The is a result of exactly solving each reduced system as described in Section˜B.2. For SPAI with three levels of infill, performing quantum multiplication requires encoding two banded matrices with 5 and 41 diagonals. Whereas, classical multiplication encodes a single banded matrix with 21 diagonals. Asymptotically, this reduces the complexity of forming $PA$ from $O(i^{2})$ to $O(i)$. See Section˜3.4.1 for a discussion of the classical pre-processing costs.
@@ 3 Preconditioning
As a test case for preconditioning, we use CFD pressure correction matrices for the flow in a lid-driven cavity taken from the open source qc-cfd  [^1] https://github.com/rolls-royce/qc-cfd/tree/main/2D-Cavity-Matrices repository. The CFD meshes range from $4\times 4$ to $64\times 64$ cells. The corresponding pressure correction matrices have dimensions ranging from $16\times 16$ to $4,096\times 4,096$ and are each extracted from the non-linear solver after 100 outer iterations.
@@
We study the impact of preconditioning on subnormalisation and subnormalised condition number, where Table˜2 gives an overview of the scenarios considered. Firstly, the preconditioners are evaluated exactly as they would be used by a classical solver (Section 3.1). Then we study the subnormalisation factors for each preconditioner (Section 3.2), required for preconditioning by quantum multiplication. In Sections 3.3 and 3.4, we study preconditioning with the quantum multiplication and classical multiplication approaches, respectively. For preconditioning by classical multiplication, only TPAI and SPAI are considered as the CLAI preconditioner produces a dense matrix that would require excessive classical computing resources. The effects of preconditioning on both subnormalisation factors and condition numbers are considered.
As mentioned above diagonal scaling is a preprocessing method used by all the preconditioners. It is included in the classical section only to show the small effect it has. The SPAI and TPAI results include infill up to three levels. All matrices are scaled to have a max norm of 1 prior to block encoding.
@@ 3.1 Classical preconditioning
Figure 5: Classical preconditioning - condition numbers for $PA$ as used by a classical algorithm. Computed using matrix-matrix multiplication. Numbers in brackets indicate levels of infill. All results are pre-scaled as described in Section˜2.2 except the ’No Scaling’ line in (a).
Figure˜5 shows the effect of the four preconditioners on the condition numbers of the CFD matrix. As discussed in Section˜2.2, diagonal scaling has little benefit by itself. There is some erratic behaviour for SPAI and TPAI with higher levels of infill on the coarser meshes which is due to the small matrix size. For larger meshes, there is an increasing benefit of more levels of infill. TPAI shows an initial improvement without infill but only a marginal improvement thereafter. There is no option to use infill with CLAI. However, it performs as well as SPAI with 3 levels of infill for the larger matrices.
@@ 3.2 Subnormalisation factors of preconditioners
[…]
From Figure˜6(a) we can see that the subnormalisation factor for $A$ is $s=3$ for all meshes. This is due to the mass conservation equation being used to construct the pressure correction matrix, where for each row the diagonal entry is the negative of the sum of the off diagonal entries. Due to 1-sided differencing at the boundaries and the diagonal pre-scaling, the largest entry along each off-diagonal is 0.5. The diagonal pre-scaling sets all entries on the main diagonal to 1.
(a) Unpreconditioned CFD Matrix AA, $s$.
@@
The first observation is the striking difference between the values of $s$ for the SPAI and TPAI preconditioners, Figure˜6(b) and Figure˜6(c), as the level of infill is increased. This is a combination of the numbers of diagonals created by the different infill strategies and the scale factor $r_{p}$. The numbers of diagonals added for each level of infill are compared in Table˜1. These are independent of the CFD mesh size except for the two smallest cases.
@@
The second observation is that for the circulant preconditioner, $s$ rises rapidly with the size of the matrix. This is due to the fact that as the mesh size increases, the cell dimensions become smaller and the lowest eigenvalue reduces. The increase in $s$ is driven entirely by the need to scale $\Lambda^{-1}$ so that $\Lambda^{-1}||_{max}=1$. There is no subnormalisation factor associated with the quantum Fourier transforms.
(a) No preconditioning, $\kappa_{s}$.
@@
Figure 7: Preconditioning by quantum multiplication - condition numbers using all preconditioners. The solid blue line with circles gives the classical condition numbers ($\kappa$) of AA without preconditioning. The dashed green line with inverted triangles gives the encoded condition numbers ($\kappa_{s}$) from $U_{A}$ without preconditioning. Numbers in brackets indicate levels of infill. All results include diagonal pre-scaling.